Status and Schedule

Data refreshes automatically every 5 minutes. Click an image to enlarge.

Recent Status Messages

Want to receive APS status notifications automatically?
Jul, 31 2014 01:51:55 Top-Up Aborted at 01:48
Top Up aborted at 01:48 due to Real time feedback. Top up restarted at 01:50.
Jul, 30 2014 11:57:10 SR beam motion from 11:44 to 11:50
There was minor SR beam orbit motion from 11:44 to 11:50 due to an issue with Realtime Feedback.
Jul, 30 2014 11:37:07 SR beam orbit motion from 11:14 to 11:20
There was a minor SR beam orbit motion from 11:14 to 11:20, and the orbit correction was not running in the horizontal plabe because of a glitch at 34 ID downstream gap readback.
Jul, 30 2014 10:04:29 Shutters enabled
On July 30, 2014, at 09:31AM. Stored beam was lost. Shutter permit was enabled at 10:00AM. There are ~100mA of stored beam, Top-Up is ongoing.
Jul, 30 2014 09:37:43 Stored beam lost
On July 30, 2014, at 09:31AM. Stored beam was lost. Investigation is ongoing. Next Fill info ~10:30AM.
Jul, 27 2014 01:26:20 Shutter permit enabled at 01:25, on Sunday, July 27th.
Shutter permit was enabled at 01:25 on Sunday, July 27th. There is 102mA of stored beam. Top-Up is ongoing for Fill #26.
Jul, 27 2014 00:55:03 Shutter Permit disabled at 00:51.
Shutter permit was disabled at 00:51, on Sunday, July 27th. Stored beam was lost. Fill #25 is over.
Jul, 26 2014 07:16:39 Shutter Permit Enabled at 07:13, on Saturday, July 26th.
Shutter permit was enabled on Saturday, July 26th at 07:13. There is 102mA of stored beam. Top-up is ongoing for fill #25.
Jul, 26 2014 06:43:00 Stored beam loss update
Stored beam lost on Saturday July 26, 2014 at 01:11. At this time Diagnostics and Controls Group personnel are at the APS continuing the investigation of the situation. Thank You, Steven LaBuda
Jul, 26 2014 05:09:08 Stored beam loss update
Stored beam lost on Saturday July 26, 2014 at 01:11. At this time Diagnostics and Controls Group personnel are at the APS investigating the situation. Michael Borland has been infromed of the extended downtime. Thank You, Steven LaBuda
Jul, 26 2014 03:08:15 Stored beam loss update
Stored beam lost on Saturday July 26, 2014 at 01:11. At this time Operations, Karen Schroder, Louis Emery, Hanh Bui (at APS), and Tony Pietryla (at APS) are investigating the situation. Thank You, Steven LaBuda
Jul, 26 2014 01:14:40 Shutter Permit disabled at 1:11.
Shutter permit was disabled on at 01:11, on Satuday, July 26th. Stored beam was lost. investigation is ongoing.
Jul, 26 2014 01:11:10 Top-Up aborted at 1:10
Top-Up aborted at 1:10, on Saturday, July 26th. Investigation is ongoing.
Jul, 26 2014 01:06:35 Top-Up Restarted
Top-Up restarted at 1:05, on Saturday, July 26th.
Jul, 26 2014 00:50:14 Top-Up Aborted due to a control law issue, at 00:47.
On Saturday, July 26, at 00:47 Top-Up was aborted due to a control law issue. Investigation is ongoing.
Jul, 24 2014 00:39:51 Top-Up resumed at 00:39hrs
Jul, 24 2014 00:36:29 Top-up aborted @ 00:33hrs due an injector issue.
Jul, 23 2014 15:01:05 Top-up interrupted
Due to an injector issue, top-up was suspended from 14:50 til 14:58.
Jul, 21 2014 11:19:59 Top-up restarted 11:18
After rebooting an IOC, top-up has been restarted at 11:18 with 96mA stored.
Jul, 21 2014 10:53:49 Top-up suspended 10:49
Due to an IOC issue, top-up was suspended at 10:49.
Jul, 21 2014 05:21:08 Top-up resumed
At 05:19, Top-up and orbit control have resumed.
Jul, 21 2014 05:15:51 Top-up aborted
At 05:11, top-up and orbit control were aborted due to an IOC issue.
Jul, 20 2014 02:33:35 Top-up Resumed
The boot of IOCB7BPM was completed. Top-up resumed at 02:28. OPS
Jul, 20 2014 02:21:56 Top-up suspended
On Sunday July 20 at 02:18, top-up has been temporarily suspended due to an IOC issue. OPS
Jul, 19 2014 19:17:52 Beam lost due to S17A:Q4 Power Supply Trip
On Saturday, July 19th, stored beam was lost due to S17a:q4 power supply trip. Shutter permit for Fill #23 was enabled @18:34 with 102mA of stored beam. Top-Up is ongoing.
Jul, 19 2014 16:29:28 Top-Up Started
The Boot of IOCS1BPM is complete. Top-Up started at 16:29.
Jul, 19 2014 16:18:18 Top-Up Aborted
On Saturday, July 19th, Top-up was aborted to boot IOCS1BPM
Jul, 19 2014 15:02:40 Top-Up Started
The Boot of IOCS3BPM and IOCS5BPM is complete. Top-Up was started at 15:02
Jul, 19 2014 14:59:16 Top-Up aborted due to IOCS3BPM and IOCS5BPM
On Saturday, July 19th at 14:56 Top-up was aborted to boot IOCS3BPM and IOCS5BPM.
Jul, 19 2014 10:20:36 Shutter permit for Fill #22 enabled @ 10:20hrs.
There is 102mA stored. Top-Up is ongoing.
Jul, 19 2014 10:00:01 Stored beam lost @ 09:45
Next stored beam @ ~10:30hrs.
Jul, 19 2014 05:12:39 Top-Up and Orbit control aborted.
Top-Up and Orbit were aborted from 04:59 to 05:11 to correct an IOC issue.
Jul, 18 2014 22:08:52 Top-Up started
The boot of IOCS7BPM is complete. On Friday July 18th 2014 at 22:08 top up was started.
Jul, 18 2014 22:03:46 On Friday July 18th 2014 at 21:57 top up was aborted due to IOCS7BPM.
On Friday July 18th 2014 at 21:57 top up was aborted to boot IOCS7BPM.
Jul, 18 2014 21:13:38 top-up started
On Friday July 18th 2014 at 21:01 top up was aborted to boot IOCS7BPM. The boot of IOCS7BPM is complete. On Friday July 18th 2014 at 21:09 top up was started.
Jul, 18 2014 21:04:31 top-up aborted due to IOCS7BPM
On Friday July 18th 2014 at 21:01 top up was aborted to boot IOCS7BPM.
Jul, 17 2014 01:22:36 Top-up / Realtime Feedback / SR Orbit Correction Resumed.
On July 16, 2014 at 01:09AM , Top-Up was aborted due to an IOC issue. at 01:10AM - Realtime Feedback and Storage Ring orbit correction were aborted. The IOC issues were resolved and at 01:19 AM , Top-Up, Realtime Feedback, and SR Orbit Correction were resumed. - Ops.
Jul, 17 2014 01:13:04 Top-Up / Realtime Feedback / SR Orbit Correction aborted.
On July 16, 2014 at 01:09AM , Top-Up was aborted due to an IOC issue. at 01:10AM - Realtime Feedback and Storage Ring orbit correction were aborted. - Ops.
Jul, 17 2014 00:08:28 Top-up / Realtime Feedback / SR Orbit Correction Resumed.
On July 16, 2014 at 11:49PM , Top-Up was aborted due to an IOC issue. at 11:50 - Realtime Feedback and Storage Ring orbit correction were aborted. The IOC issues were resolved and at 12:05 AM , Top-Up, Realtime Feedback, and SR Orbit Correction were resumed. - Ops.
Jul, 16 2014 23:54:10 Top-Up / Realtime Feedback / SR Orbit Correction aborted.
On July 16, 2014 at 11:49PM , Top-Up was aborted due to an IOC issue. at 11:50 - Realtime Feedback and Storage Ring orbit correction were aborted. - Ops.
Jul, 11 2014 14:21:17 Shutter permit was enabled on Friday, July 11th at 14:06.
Shutter permit was enabled on Friday, July 11th at 14:06. Fill #19 has begun. Top-up is ongoing with ~102mA of stored beam.
Jul, 11 2014 13:38:54 On Thursday, July 11th, at 13:26 Stored Beam was lost due to water flow fault.
On Thursday, July 11th, at 13:26 Stored Beam was lost due to water flow fault.
Jul, 09 2014 07:43:32 Fill on fill with Shutters Open was completed on Wednesday, July 9th @7:41. ~102mA of Stored Beam
Fill on fill with shutters open was completed on Wednesday, July 9th @7:41. ~102mA of Stored Beam.
Jul, 09 2014 07:12:59 FW: CORRECTION: Top-up Has Been restarted on Wednesday, July 9th at 07:07. In-Reply-To: References:
________________________________ From: asdops To: asdops Subject: CORRECTION: Top-up Has Been restarted on Wednesday, July 9th at 07:07. CORRECTION: Top-up has been restarted on Wednesday, July 9th at 07:12. There will be a fill on fill with shutters open at 07:40.
Jul, 09 2014 07:10:44 Top-Up Has Been Restarted on Thursday, July 7th at 07:07.
Top-Up has been restarted on Thursday, July 7th at 07:07. There will be a fill on fill with shutters open at 07:40.
Jul, 09 2014 06:19:13 Top-up scheduled 07:15
Booster dipole repairs are complete, however the power supply is being monitored for an hour before it is released for top-up operations at ~07:15.
Jul, 09 2014 05:49:53 Booster dipole update
A broken wire for the Booster dipole Choke temperature interlock circuit needs to be re-terminated, the work is estimated to take ~30-45minutes.
Jul, 09 2014 05:05:17 103mA, no top-up
At 04:45, the fill-on-fill to 103mA was completed and top-up ran for another 5 minutes before the Booster dipole tripped again. Estimated time of repair is unknown, updates will be forthcoming.
Jul, 09 2014 04:20:40 Fill-on-fill at 04:40
A fill-on-fill with shutters open has been scheduled for 04:40. Top-up is not available at the moment.
Jul, 09 2014 04:05:40 Coupling change cancelled
The SR coupling change planned for 03:41 has been cancelled. Repairs to the injector are expected to be complete before a coupling change is necessary.
Jul, 09 2014 03:36:08 Booster dipole tripped again
The Booster dipole tripped again after repairs were completed, top-up only ran for 5 minutes (03:04-03:09). In 5 minutes, we will be making an adjustment to SR coupling in order to increase beam lifetime.
Jul, 09 2014 02:36:50 Booster dipole repairs ongoing
Repairs to the Booster dipole are ongoing, no estimate for completion is available. A fill-on-fill to 102mA will be performed 30minutes after the power supply gets fixed.
Jul, 09 2014 00:47:34 Injector repairs
Technicians are en route to the APS to repair the Booster dipole power supply, their ETA is ~1 hour. Additional time will be needed to perform the work.
Jul, 09 2014 00:20:13 Top-up aborted at 00:12
Due to an injector issue, top-up was aborted at 00:12. There are no estimates for repair at this time, updates will be forthcoming.
Jul, 04 2014 06:36:42 Storage Ring Orbit Correction
Workstation orbit correction has been started on both planes at 06:19am.
Jul, 04 2014 06:01:21 Stored Beam
At 05:34 we injected into the timing bunch. At 05:51 we completed the fill with 102mA of stored beam and started top-up. At this time orbit correction is not working due to the network issue.
Jul, 04 2014 05:17:57 Control systems still not functioning, workaround underway
Alternative methods to restore the control systems to operations are still under way. As of now, there is no estimated time for a refill. We are attempting to restore the ability to inject into the Storage Ring first.
Jul, 04 2014 02:28:50 Control systems still not functioning, workaround underway
Alternative methods to restore the control systems to operations are still under way. As of now, there is no estimated time for a refill. We are attempting to restore the ability to inject into the Storage Ring first. Restoration of orbit control will take much longer, if possible at all with the current workaround.
Jul, 04 2014 00:58:42 Storage ring injection still unavailable.
The restoration of the Linux file system has not proceeded as expected. Alternative methods to restore the control systems to operations are being attempted. As of now, there is no estimated time for a refill. We dropped below 50mA at approximately 23:00. We will send out an update when we determine if any of these alternative methods will be successful or not.
Jul, 03 2014 21:31:30 No Storage Ring injection due to network problems. - update.
On July 3, 2014 at 15:37, stored beam dropped below 100mA; At 5:54PM, top up was aborted. IT personnel has successfully recovered the servers. Currently, the linux filesystem is being restored; it is expected to take 2-3 hours to finish recovery. Next-Fill update ETA ~11:00PM. - Ops
Jul, 03 2014 19:58:31 No Storage Ring injection due to network problems. - update.
On July 3, 2014 at 15:37, stored beam dropped below 100mA; At 5:54PM, top up was aborted. The server has been recovered. However, our OAG applications directory is still being loaded. Next fill update ETA ~1 hour. -Ops
Jul, 03 2014 19:02:37 No Storage Ring injection due to network problems. - update.
On July 3, 2014 at 15:37, stored beam dropped below 100mA; At 5:54PM, top up was aborted. Server repairs are ongoing. Next fill update ETA ~1 hour.
Jul, 03 2014 17:59:48 Top-Up aborted.
On July 3, 2014 at 5:54PM, top up was aborted. network investigation and repairs are ongoing. -Ops
Jul, 03 2014 17:02:17 No Storage Ring injection due to network problems - update.
On July 3, 2014 at 15:37, stored beam dropped below 100mA. Top-Up is still running but there is no beam being injected. The network problems are ongoing, IT personnel are working with the vendors to resolve the issue. Due to multiple MCR Workstations being down, Storage Ring orbit correction is not running at this time and cannot be restarted. Next fill update ETA ~1 hour. -Ops
Jul, 03 2014 17:02:08 No Storage Ring injection due to network problems - update.
On July 3, 2014 at 15:37, stored beam dropped below 100mA. Top-Up is still running but there is no beam being injected. The network problems are ongoing, IT personnel are working with the vendors to resolve the issue. Due to multiple MCR Workstations being down, Storage Ring orbit correction is not running at this time and cannot be restarted. Next fill update ETA ~1 hour. -Ops
Jul, 03 2014 15:52:49 No Storage Ring injection due to network issues.
On July 3, 2014 at 3:37PM, stored beam dropped below 100mA. Top-Up is still running but there is no beam being injected into the Storage Ring. Investigation of ongoing Network issues is ongoing. -Ops.